# LLMs Cannot Reliably Judge (Yet?): A Comprehensive Assessment on the Robustness of LLM-as-a-Judge

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable intelligence across various tasks, which has inspired the development and widespread adoption of LLM-as-a-Judge systems for automated model testing, such as red teaming and benchmarking. However, these systems are susceptible to adversarial attacks that can manipulate evaluation outcomes, raising concerns about their robustness, hence trustworthiness. Existing evaluation methods adopted by LLM-based judges are often piecemeal and lack a unified framework for comprehensive assessment. Furthermore, the prompt template and model selections for improving judge robustness have been rarely explored, and their performance in real-world settings remains largely unverified. To address these gaps, we introduce RobustJudge, a fully automated and scalable framework designed to systematically evaluate the robustness of LLM-as-a-Judge systems. RobustJudge investigates the impact of attack methods and defense strategies (RQ1), explores the influence of prompt template and model selection (RQ2), and assesses the robustness of real-world LLM-as-a-Judge applications (RQ3). Our main findings are that 1) LLM-as-a-Judge systems are still vulnerable to a range of adversarial attacks, including Combined Attack and PAIR, while defense mechanisms such as Re-tokenization and LLM-based Detectors offer improved protection. 2) Robustness is highly sensitive to the choice of prompt template and judge models. Our proposed prompt template optimization method can improve robustness, and JudgeLM-13B demonstrates strong performance as a robust open-source judge. 3) Applying RobustJudge to Alibaba’s PAI platform reveals previously unreported vulnerabilities. The source code of RobustJudge is provided at https://github.com/S3IC-Lab/RobustJudge.
1 Introduction
@@
Large Language Models (LLMs), such as OpenAI’s GPT-4o [1], Google’s Gemma2 [2], Meta’s Llama 3 [3], and QwenLM’s Qwen2.5 [4], have achieved remarkable proficiency across a wide range of tasks. Built on extensive training data and Transformer architectures, these models demonstrate advanced capabilities in general-purpose intelligence, including natural language understanding, text generation, and complex problem-solving.
@@
To leverage these capabilities while minimizing human effort and mitigating human biases in LLM evaluation, the concept of LLM-as-a-Judge has been introduced. LLM-as-a-Judge [5] aims to automate the assessment of LLM-generated content, providing a scalable and objective alternative to human evaluation. This approach has gained widespread adoption, becoming a defacto evaluation method for assessing LLM performance across various domains, including software engineering[6], domain-specific knowledge assessment[7], and mathematical reasoning[8].
@@
The initial success of LLM-as-a-Judge can be attributed to its strong agreement with human preferences on standard benchmarks. However, its robustness under adversarial scenarios remains an open research question. Recent studies [9, 10, 11, 12] have revealed that LLM-as-a-Judge systems are intrinsically vulnerable to various forms of adversarial attacks, which can manipulate evaluation outcomes without introducing easily detectable anomalies. These findings raise significant concerns about the reliability of LLM-as-a-Judge systems.
@@
Given the increasing reliance on LLM-as-a-Judge platforms for LLM evaluation, such as AlpacaEval [13], Chatbot Arena [5], and MT-Bench [5], ensuring their robustness has become a pressing research priority. However, existing assessment methods for these systems are fragmented, lacking a unified, systematic, and automated framework for comprehensive evaluations. Furthermore, the optimal prompt configuration and model selection for different tasks and evaluation protocols are underexplored, with limited guidance on best practices. Finally, while real-world LLM-as-a-Judge applications have recently demonstrated promising effectiveness in content evaluation, their robustness in adversarial settings remains unverified.
@@
Our Work. To address these challenges, we introduce RobustJudge, a fully automated and scalable framework to evaluate the robustness of LLM-as-a-Judge systems. Our framework systematically assesses these systems by exploring three core research questions:
@@
  • RQ1: What impact do different adversarial attacks and defense methods have on the LLM-based judges?
@@
  • RQ2: How do the prompt templates and model choices affect the robustness of LLM-based judges?
@@
  • RQ3: What vulnerabilities exist in black-box real-world deployments of LLM-based judges, as revealed by our empirical evaluations?
@@
To answer RQ1, we conduct a comprehensive evaluation of LLM-based judges against 15 adversarial attack techniques and 7 defense strategies. Our analysis provides extensive comparisons of these techniques, revealing critical insights into their relative strengths and weaknesses. The results indicate that many attack methods, such as Fake Reasoning [12], Combined Attack [14], Empty Attack and AdvEval [10], PAIR [15], TAP [16] consistently achieve high attack success rate on multiple tasks and models, highlighting vulnerabilities in LLM-based judges. Conversely, defense strategies such as Retokenization [17] and Naive LLM-based Detector [18], demonstrate notable effectiveness in mitigating these attacks. Our evaluation provides a comprehensive view of the current landscape of attack and defense techniques, and a clear guidance on safeguarding LLM-as-a-Judge systems against adversarial manipulation.
@@
To address RQ2, we analyze the effect of adversarial attacks under different judge prompt templates and judge model selection. Our analysis reveals that the robustness of LLM-as-a-Judge systems is highly sensitive to both factors. Specifically, while all the evaluated prompt templates are clear and well-structured, they present different levels of robustness to the LLM-as-a-Judge system. To mitigate this issue, we propose a prompt template optimization method aimed at identifying configurations with improved robustness. Our optimized template consistently outperforms existing templates in robustness against multiple attacks. Additionally, our evaluation on various judge models reveals that JudgeLM-13B exhibits substantially stronger robustness against adversarial inputs, performing comparably to the widely adopted GPT-4o. This suggests that JudgeLM-13B is a strong open-source alternative as an LLM judge. These findings emphasize the importance of prompt template design and judgment fine-tuning in enhancing the robustness of LLM-as-a-Judge systems against adversarial attacks.
@@
To study RQ3, we investigate the robustness of a real-world LLM-based judge system, i.e., Alibaba’s PAI platform. Our evaluation leverages the adversarial examples generated by RobustJudge and is conducted via the public API provided by the platform. The results show that conventional adversarial attacks were ineffective against PAI-Judge platform, suggesting a strong level of robustness. However, we identify a critical loophole using a composite attack, which combines PAIR-optimized adversarial inputs with long-suffix manipulations. This strategy successfully bypasses the platform’s defenses and alters the model’s evaluation outcome. These findings indicate that RobustJudge is practical in identifying hidden vulnerabilities and guiding the development of more robust LLM-as-a-Judge systems.
@@
  • We develop RobustJudge, the first fully automated and scalable framework designed for extensive robustness evaluation of LLM-as-a-Judge systems.
@@
  • RobustJudge evaluate 15 adversarial attacks and validates 7 judge defense strategies. Our analysis reveals vulnerabilities of LLM judges on various tasks, consistency of judgment across evaluation protocols, and high attack success rates from methods such as Combined Attack and PAIR. While defenses like Re-tokenization and LLM-based Detectors show promise, their effectiveness often comes with notable trade-offs in accuracy or usability.
@@
  • Utilizing RobustJudge, we conduct an in-depth investigation on the configuration of LLM-as-a-Judge systems, focusing on the prompt templates and judge model selection. Our extensive analysis identifies the most reliable configuration against adversarial attacks, offering actionable guidance for enhancing system robustness.
@@
  • We evaluate a real-word LLM judge system Alibaba’s PAI platform using RobustJudge, and identify hidden vulnerability of the platform. We share our findings with the PAI team to support the development of more robust LLM-as-a-Judge systems.
2 Background and Related Work
2.1 LLM-as-a-Judge
Large language models (LLMs) have exhibited state-of-the-art proficiency in understanding and generating human-like text. Despite their impressive capabilities, traditional reference-based metrics such as ROUGE [19] and BLEU [20] are limited in capturing the nuanced quality of their outputs. To overcome these limitations, Zheng et al. proposed the LLM-as-a-Judge paradigm [5], which employs LLMs themselves as evaluators to assess the quality of model outputs. This paradigm has rapidly gained popularity, leading to the development of a variety of fine-tuned judge models tailored for different evaluation settings [21].
For example, Wang et al. developed PandaLM [22], which avoids costly API calls and mitigates privacy risks. Zhu et al. proposed JudgeLM [23], enhancing accuracy via swap augmentation and reference support. Kim et al. released Prometheus 2 [24], an open-source evaluator achieving top alignment with human and GPT-4 judgments. Beyond general text, the judge paradigm has been adapted to specialized domains, including software engineering [6], machine translation [25], legal reasoning [7], and mathematical problem solving [8]. Notably, Zhuge et al. introduced Agent-as-a-Judge [26], leveraging autonomous agents as evaluators.
@@ 2.1 LLM-as-a-Judge
Several benchmarks have also been developed to evaluate the judge’s performance. Early efforts like LLMEval [27], MTBench [5], and FairEval [28] focus on consistency with human judgment, though they may favor stylistic fluency over factual correctness. LLMBar [29] improves evaluation rigor by introducing ground-truth preference labels and stricter instruction adherence. JudgeBench [8] further advances the field by emphasizing objective correctness in challenging domains such as knowledge, reasoning, mathematics, and programming.
However, while these benchmarks assess overall judging ability, they do not examine evaluator safety or robustness. In this work, we introduce the first benchmark designed specifically to evaluate the safety risks and reliability of LLM-based evaluators.
Figure 1: LLM-as-a-judge pipeline.
2.2 Attacks against LLM-as-a-Judge
We categorize attacks on LLM-as-a-Judge into two main types: heuristic-based and optimization-based.
Heuristic-based Attacks. Heuristic-based attacks refer to adversarial strategies that are manually crafted or rule-based, exploiting model behaviors without relying on gradient information or automated optimization. Typically, these attacks leverage prompt engineering and input manipulation to coerce the model into unintended behaviors, and can be broadly categorized into prompt injection attacks and adversarial example attacks.
In prompt injection attacks, a basic approach is to concatenate the benign input with an adversarial sequence [30, 31, 32]. Alternatively, inserting special tokens or delimiters can activate hidden or unauthorized behaviors [31].
Optimization-based Attacks. Optimization-based attacks leverage gradient signals or structured search procedures to systematically craft inputs that achieve specific adversarial objectives. In this section, we focus on both jailbreak attacks and adversarial strategies.
Beyond prompt injection, Chen et al. [10] introduced AdvEval, a black-box adversarial framework that generates targeted examples against NLG evaluators using feedback from powerful LLMs. By iteratively refining inputs based on discrepancies between human and model judgments, AdvEval substantially degrades the performance of multiple evaluation metrics across diverse tasks and datasets.
Jailbreaking involves designing inputs that circumvent an LLM’s built-in safeguards, coercing it into producing content that violates its usual policy constraints, such as harmful or unethical outputs [33, 16, 34, 15]. However, conventional jailbreak techniques cannot be directly applied to LLM-as-a-Judge, since the evaluation setting does not reduce to simple suffix optimization. To bridge this gap, we propose a novel adaptation framework that integrates existing jailbreak methodologies into the LLM-as-a-Judge context, enabling targeted manipulation of the judge’s scoring behavior.
Recent studies have directly targeted LLM-based evaluators via adversarial manipulations. Zheng et al. [11] demonstrated that a null model—always outputting a constant, irrelevant response—can still secure high win rates on benchmarks like AlpacaEval 2.0 and MT-Bench, exposing fragility of benchmark-driven evaluation. Raina et al. [9] revealed that appending short, task-agnostic adversarial phrases can dramatically inflate scores in absolute scoring setups, underscoring the lack of robustness in LLM judges. Shi et al. [12] introduced JudgeDeceiver, a gradient-based prompt injection method that optimizes adversarial sequences to reliably mislead LLM evaluators across tasks—ranging from LLM-powered search to RLAIF—surpassing the effectiveness of manual prompt attacks.
In this paper, we integrate a wide range of adversarial strategies into a unified framework to systematically evaluate their effectiveness against LLM-as-a-Judge systems.
2.3 Defenses for LLM-as-a-Judge
We categorize defense strategies against attacks on LLM-as-a-Judge into two main types: prevention-based and detection-based approaches.
Prevention-based Defenses. Prevention-based defenses seek to block prompt injection and adversarial manipulation by preprocessing both input instructions and candidate outputs before evaluation. For instruction preprocessing, methods have been developed to refocus the model on the genuine task and neutralize injected prompts. For instance, Sandwich [35] inserts an instruction-reinforcing prompt after the input to maintain task integrity. For candidate response processing, Jain et al. [17] employed paraphrasing and retokenization to obscure adversarial triggers, and Liu et al. [14] further adapted these techniques to counter more sophisticated injections. Additionally, Li et al. [36] used LLMs to mask and reconstruct the input, disrupting adversarial continuity and substantially lowering attack success rates. Nevertheless, these generation-focused strategies often prove inadequate for LLM-as-a-Judge, where attackers target score distortions rather than harmful content generation.
Detection-based Defenses. Detection-based defenses focus on identifying potential attacks at either the input or output stages of evaluation. On the input side, Jain et al. [17] introduced a self-perplexity filter to detect prompt injection by identifying anomalous perplexity signals. Follow-up work [37] enhanced this approach by combining perplexity and token-length features to train a classifier for detecting adversarial prompts. While these input-level methods offer some protection, they often struggle to generalize to more sophisticated adversarial strategies—such as optimization-based prompt injections targeting judge LLMs [12] or adversarial input crafting techniques [9, 10] that directly manipulate score assignment or response ranking.
On the output side, Helbing et al. [38] proposed using an auxiliary LLM to verify whether generated outputs exhibit jailbreak behaviors. Similarly, sandwich-based detection [35] evaluates whether outputs remain aligned with predefined system objectives. However, such output-level defenses are generally less effective in the judge setting, where both benign and adversarial candidate responses may appear plausible. In these cases, the attack’s goal is not to generate visibly harmful content but to subtly distort scoring or ranking decisions.
In this paper, we extend existing defenses to a broad range of LLM-as-a-Judge tasks and evaluate their robustness against adversarial attacks.
3 Threat Model
We identify two principal roles: the attacker, who crafts attacks against LLM-as-a-Judge, and the defender, who safeguards the evaluation’s reliability and safety.
3.1 Attacker
@@
We categorize adversaries against LLM-as-a-Judge systems into two classes: heuristic-based and optimization-based attackers. Heuristic-based adversaries, lacking direct access to the evaluation metric $\mathit{m}$, craft adversarial inputs using indirect cues, such as score feedback [10] or prompt manipulation techniques [31, 30, 32, 39, 40, 14]. In contrast, optimization-based adversaries have full visibility into the evaluation prompt and scoring rules, enabling white-box optimization to directly maximize their evaluation outcome [11, 12, 9].
Attacker’s Capabilities. Regardless of attacker type, we assume the attacker can manipulate the response submitted the LLM judge for evaluation. In particular, the attacker is allowed to inject arbitrary instructions or sequences into the malicious response. For example, the attacker may insert additional content at any position within the response to increase its score or the likelihood of being selected. However, attackers are not allowed to modify the judge’s instruction prompt or internal model parameters, which are fixed and controlled by the host institution. Besides, attackers do not have access to the competing candidate’s response $r$ in pairwise evaluation settings.
Attacker’s Goal. The attacker’s goal is to manipulate the LLM-as-a-Judge pipeline by crafting responses that skew evaluations in their favor.
In the scoring scenario, the attacker generates a malicious response $r_{m}$ designed to artificially maximize the probability of the top score without genuinely meeting the criteria. Formally, it solves:
@@
where $w_{k}$ is the token for the highest evaluation score.
In the pairwise comparison scenario, the attacker forges $r_{m}$ to ensure it is chosen over a benign response $r_{b}$, i.e.:
@@
where $w_{m}$ indicates selection of the malicious response.
3.2 Defender

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional capabilities across diverse tasks, driving the development and widespread adoption of LLM-as-a-Judge systems for automated evaluation, including red teaming and benchmarking. However, these systems are susceptible to adversarial attacks that can manipulate evaluation outcomes, raising critical concerns about their robustness and trustworthiness. Existing evaluation methods for LLM-based judges are often fragmented and lack a unified framework for comprehensive robustness assessment. Furthermore, the impact of prompt template design and model selection on judge robustness has rarely been explored, and their performance in real-world deployments remains largely unverified. To address these gaps, we introduce RobustJudge, a fully automated and scalable framework designed to systematically evaluate the robustness of LLM-as-a-Judge systems. Specifically, RobustJudge investigates the effectiveness of 15 attack methods and 7 defense strategies across 12 models (RQ1), examines the impact of prompt template design and model selection (RQ2), and evaluates the security of real-world deployments (RQ3). Our study yields three key findings: (1) LLM-as-a-Judge systems are highly vulnerable to attacks such as PAIR and combined attacks, while defense mechanisms such as re-tokenization and LLM-based detectors can provide enhanced protection; (2) robustness varies substantially across prompt templates (up to 40%); (3) deploying RobustJudge on Alibaba’s PAI platform uncovers previously undiscovered vulnerabilities. These results offer practical insights for building trustworthy LLM-as-a-Judge systems.
I Introduction
@@
Large Language Models (LLMs), such as OpenAI’s GPT-4o 2024gpt4 , Google’s Gemma2 gemma2 , Meta’s Llama 3 llama3 , and QwenLM’s Qwen2.5 2025qwen , have achieved remarkable proficiency across a wide range of tasks. Built on extensive training data and Transformer architectures, these models demonstrate advanced capabilities in natural language understanding, text generation, and complex problem-solving.
@@
To leverage these capabilities while minimizing human effort and mitigating human biases in LLM evaluation, the concept of LLM-as-a-Judge has been introduced. LLM-as-a-Judge judgeFirst aims to automate the assessment of LLM-generated content, providing a scalable and objective alternative to human evaluation. This approach has gained widespread adoption, becoming an extensive evaluation method for assessing LLM performance across various domains, including software engineering codeJudge , domain-specific knowledge assessment legalJudge , and mathematical reasoning judgebench .
Figure 1: An overview of our judge prompt’s structure, composed of Component Design, Evaluation Process, and Evaluation Metrics. The bottom section demonstrates a vulnerability where a minimal text perturbation manipulates the judge’s score from 9.3 to 10.0.
@@
The initial success of LLM-as-a-Judge can be attributed to its strong agreement with human preferences on standard benchmarks. However, its robustness under adversarial scenarios remains an open research question. Recent studies op2 ; adv ; cheating ; judgeDeceiver have revealed that LLM-as-a-Judge systems are inherently vulnerable to various forms of adversarial attacks, which can manipulate evaluation outcomes through subtle perturbations. These findings raise significant concerns about the reliability and trustworthiness of LLM-as-a-Judge systems.
@@
Given the increasing reliance on LLM-as-a-Judge platforms for LLM evaluation, such as AlpacaEval alpacaEval , Chatbot Arena judgeFirst , and MT-Bench judgeFirst , ensuring their robustness has become a pressing research priority. However, existing assessment methods for these systems are fragmented, lacking a unified, systematic, and automated framework for comprehensive evaluation. Furthermore, the optimal prompt configuration and model selection for different tasks and evaluation protocols are insufficiently explored, with limited guidance on best practices. Finally, while real-world LLM-as-a-Judge applications have recently demonstrated promising effectiveness in content evaluation, their robustness in adversarial settings remains largely unverified.
@@
Our Work. To address these challenges, we introduce RobustJudge, a fully automated and scalable framework designed to evaluate the robustness of LLM-as-a-Judge systems systematically. Our framework assesses these systems by exploring three core research questions:
@@
  • RQ1: What impact do different adversarial attack methods and defense strategies have on LLM-based judges?
@@
  • RQ2: How do prompt template design and model selection affect the robustness of LLM-based judges?
@@
  • RQ3: What vulnerabilities exist in black-box real-world deployments of LLM-based judges, as revealed through empirical evaluation?
@@
To answer RQ1, we conduct a comprehensive evaluation of LLM-based judges against 15 adversarial attack techniques and 7 defense strategies. Our analysis provides extensive comparisons of these techniques, revealing critical insights into their relative strengths and weaknesses. The results indicate that many attack methods, such as Fake Reasoning judgeDeceiver and Combined Attacks prompt_bench , consistently achieve high attack success rates across multiple tasks and models, highlighting vulnerabilities in LLM-based judges. Conversely, defense strategies such as re-tokenization baseDefense and naive LLM-based detectors naive_llm_defense demonstrate notable effectiveness in mitigating these attacks. Our evaluation provides a comprehensive view of the current landscape of attack and defense techniques, offering clear guidance on safeguarding LLM-as-a-Judge systems against adversarial manipulation.
@@
To address RQ2, we analyse the effects of adversarial attacks across different judge prompt templates and model selections. Our analysis reveals that the robustness of LLM-as-a-Judge systems is highly sensitive to both factors. Specifically, while all evaluated prompt templates are clear and well-structured, they exhibit varying levels of robustness. To mitigate this issue, we propose a prompt template optimization method designed to identify configurations with enhanced robustness. Our optimized template consistently outperforms existing templates in robustness against multiple attacks. Additionally, our evaluation of various judge models reveals that JudgeLM-13B exhibits substantially stronger robustness against adversarial inputs, performing comparably to the widely adopted GPT-4o  [^1] We use GPT-4o-1120 in this paper.. The results suggest that JudgeLM-13B is a strong open-source alternative for LLM-based judging. These findings emphasize the importance of prompt template design and fine-tuning in enhancing the robustness of LLM-as-a-Judge systems against adversarial attacks.
@@
To study RQ3, we investigate the robustness of a real-world LLM-based judge system deployed on Alibaba’s PAI platform. Our evaluation leverages the adversarial examples generated by RobustJudge and is conducted via the platform’s public API. The results indicate that conventional adversarial attacks are ineffective against the PAI-Judge platform, suggesting a robust baseline level of security. However, we identify a critical vulnerability that can be exploited through a composite attack strategy combining PAIR-optimized adversarial inputs with long-suffix manipulations. This strategy successfully bypasses the platform’s defenses and alters the model’s evaluation outcomes. These findings demonstrate that RobustJudge is effective in identifying hidden vulnerabilities and can guide the development of more robust LLM-as-a-Judge systems.
@@
  • We develop RobustJudge, the first fully automated and scalable framework designed for comprehensive robustness evaluation of LLM-as-a-Judge systems.
@@
  • RobustJudge evaluates 15 adversarial attack methods and validates 7 defense strategies. Our analysis reveals that LLM-based judges exhibit significant vulnerabilities across various tasks, with attacks such as combined attacks and PAIR achieving high success rates. While defenses like re-tokenization and LLM-based detectors show promise, their effectiveness often involves trade-offs in accuracy or computational overhead.
@@
  • We conduct an in-depth investigation into the configuration of LLM-as-a-Judge systems, focusing on prompt template design and judge model selection. Our extensive analysis identifies the most robust configurations against adversarial attacks, offering actionable guidance for enhancing system robustness.
@@
  • We evaluate a real-world LLM-based judge system deployed on Alibaba’s PAI platform using RobustJudge and identify previously unreported vulnerabilities. We share our findings with the PAI team to support the development of more robust LLM-as-a-Judge systems.
II Related Work
II-A LLM-as-a-Judge
Large language models (LLMs) have demonstrated state-of-the-art performance in understanding and generating human-like text. However, traditional reference-based metrics such as ROUGE rouge and BLEU bleu are insufficient for capturing the nuanced quality of LLM-generated outputs, especially in tasks involving open-ended or multi-perspective generation.
To address this limitation, the LLM-as-a-Judge paradigm judgeFirst was introduced, which employs LLMs themselves as evaluators to assess the quality of model responses. This approach has been widely adopted in recent years, leading to the development of judge models fine-tuned for diverse evaluation settings autoj .
Several open-source judge models have been proposed. For instance, PandaLM pandalm reduces reliance on proprietary APIs and mitigates privacy concerns; JudgeLM judgelm improves evaluation fidelity through swap augmentation and reference assistance; and Prometheus 2 prometheus2 achieves strong alignment with both human and GPT-4 assessments. The paradigm has also been extended to domain-specific evaluations, including code generation codeJudge , machine translation translationJudge , legal reasoning legalJudge , and mathematical problem solving judgebench . More recently, Agent-as-a-Judge agentjudge explores the use of autonomous agents as evaluators to enable interactive, multi-turn assessments.
II-B Benchmarks and Datasets
@@
Several benchmarks have been developed to evaluate the performance of LLM-based judges. Early efforts such as LLMEval llmeval , MT-Bench judgeFirst , and FairEval faireval primarily focus on alignment with human preferences, often emphasizing stylistic fluency over factual accuracy. LLMBar llmbar enhances evaluation rigor by incorporating ground-truth preference labels and enforcing stricter instruction compliance. More recently, JudgeBench judgebench advances the field by targeting objective correctness across challenging domains, including factual knowledge, logical reasoning, mathematics, and code generation.
While these benchmarks provide valuable insights into overall evaluation quality, they overlook a critical dimension: the robustness and security of the evaluators themselves. In this work, we address this gap by proposing the first benchmark explicitly designed to assess the adversarial robustness and reliability of LLM-based judges under adversarial manipulation.
III Threat Model
III-A Threat Scenario
We focus on content-author attacks, where the adversary is an authorized content submitter exploiting the evaluation system. Unlike transmission-layer attacks (e.g., man-in-the-middle), the attacker embeds adversarial perturbations directly within their submitted content to manipulate evaluation outcomes. The attack occurs at the semantic understanding level rather than the network layer.
This threat model is realistic in numerous high-stakes scenarios:
  • Academic assessment: Students embed adversarial text in exam answers to inflate automated grading scores.
  • Model benchmarking: Researchers inject adversarial prompts into submissions to manipulate leaderboard rankings.
  • Code competitions: Participants embed adversarial comments in code to bias automated review systems.
  • Content moderation: Users insert adversarial instructions to bypass quality assessment filters.
Standard security mechanisms (e.g., digital signatures, encrypted channels) cannot defend against such attacks because: (i) the adversarial content is part of the legitimate submission, (ii) integrity verification passes as the content authentically originates from the authorized submitter, and (iii) the manipulation targets semantic understanding rather than data transmission.
III-B Attacker
@@
We categorize attackers into two classes: heuristic-based and optimization-based attackers. Heuristic-based adversaries, without access to the evaluation metric $\mathit{m}$, craft adversarial inputs using indirect cues such as score feedback adv or prompt manipulation techniques injection ; naive_2 ; context_ignore_1 ; context_ignore_2 ; fake_completion ; prompt_bench . In contrast, optimization-based adversaries have access to the evaluation prompt and scoring rules, enabling them to directly maximize evaluation outcomes through white-box optimization cheating ; judgeDeceiver ; op2 .
Attacker Capabilities. The attacker can manipulate their own response $r_{m}$ submitted for evaluation, including injecting arbitrary instructions. However, the attacker cannot: (i) modify the judge’s instruction prompt or model parameters, which are controlled by the evaluation platform; (ii) intercept or tamper with data during transmission; or (iii) access competing candidates’ responses $r$ in pairwise evaluation settings.
Attacker Goals. The attacker aims to manipulate evaluation outcomes by crafting a malicious response $r_{m}$ that receives inflated scores despite not meeting the quality requirements. In the scoring scenario, the attacker maximizes the probability of receiving the highest score:
@@
where $w_{k}$ represents the token for the highest evaluation score. In the pairwise scenario, the attacker ensures their malicious response $r_{m}$ is chosen over a benign response $r_{b}$:
@@
where $w_{m}$ indicates the selection of the malicious response.
III-C Defender